OpenStudy (anonymous):

Expand the expression and combine the like terms: -(-4-a)-4(-4a-7)

OpenStudy (anonymous):

release brackets first

OpenStudy (anonymous):

4+a-(-16a-28)=4+a+16a+28=17a+32

Nnesha (nnesha):

first you have to distribute brackets first one by -1 and second one by -4

OpenStudy (anonymous):

I have the answer now.

Nnesha (nnesha):

 its time to have explanation :) -4 times -1 -1 times -a